# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ 3.6 Asymptotic uncertainties from finite-size samples
The bias term generated by statistical fluctuations of the gradient vector and Jacobian matrix could be estimated in principle from eq. (3.2) by computing the covariance matrices of ${\bf b}$ and of the column vectors of ${{\bf A}}$, for example by bootstrap resampling, and using their nominal MC predictions $\hat{\bf b}$ and $\hat{\bf A}$ in place of their true, unknown values. The asymptotic Hessian uncertainty $\bar{\sigma}_{\rm H}={S}^{-1/2}$, i.e. the Hessian uncertainty in the limit of infinite MC statistics, could be determined from eq. (3.2) after subtracting the bias contribution.
@@
For example, for the toy model of section 2, the mean Hessian uncertainty for $N=2\times 10^{6}$, $n=200$, and $\alpha=0.03$ is found to be $\hat{\sigma}_{{\rm H}}=0.0572$ ($0.0447$) when $k=1$ ($k=0.5$), from which the asymptotic uncertainty can be predicted to be $0.0956$ from eq. (57), in good agreement with the true value of $0.0950$.
Equation (57) assumes that the likelihood does not contain any explicit term to account for the statistical uncertainties on the MC templates. It can be easily extended to the case where the MC statistical uncertainties are treated à la Barlow-Beeston lite by exploiting the Gauss + MC stat. approximation (see sec. 2.1.2), thus modifying the scaling of ${\bf b}$ via an additional multiplicative factor $\left(1+\frac{1}{k}\right)^{-{1}/{2}}$, yielding:
Equations (57) and (58) could be improved by taking the expectation values of $\hat{\sigma}_{{\rm H}}$ over several bootstrap resamplings of the MC sample.
@@
3.6.1 Lite vs full Barlow-Beeston
@@
Suppose that $\bar{\sigma}_{\rm H}$ calculated from eq. (56) is in good agreement with $\hat{\sigma}_{\rm H}$, or, equivalently, that the Hessian uncertainty is stable when the size of the MC sample is reduced. One might still want to ensure that the lite version of the Barlow-Beeston method represents an adequate approximation for a given problem. For this purpose, we can use the formalism developed in the previous section to elaborate the following analysis. We start by noticing that the MLE estimator $(\hat{\boldsymbol{\theta}},\hat{\mu})$ can be seen as a function of ${\bf B}$, ${\bf f}_{0}$, and ${\bf y}$, which are all random variables of known variance. By applying the chain rule for differentiation, the gradient of $(\hat{\boldsymbol{\theta}},\hat{\mu})$ with respect to $B_{ik}$ can be computed as
$={\bf C}\left(\boldsymbol{\Pi}^{T}_{ik}\left({\bf d}-{\bf B}\begin{pmatrix}\hat{\boldsymbol{\theta}}\\ \hat{\mu}\end{pmatrix}\right)-{\bf B}^{T}\boldsymbol{\Pi}_{ik}\begin{pmatrix}\hat{\boldsymbol{\theta}}\\ \hat{\mu}\end{pmatrix}\right),$ (59)
where $\boldsymbol{\Pi}_{ik}={\bf e}_{i}{\bf e}_{k}^{T}$, with ${\bf e}_{l}$ the $l^{\rm th}$ vector of the canonical basis. Calculating the gradient with respect to ${\bf f}_{0}$ is straightforward. If the MC predictions for $B_{ik}$ and $f_{0,i}$ can be treated as independent, Gaussian-distributed variables with standard deviations $\Delta B_{ik}$ and $\Delta f_{0,i}$, the contribution to the variance of the MLE estimator $\hat{\mu}$ induced by fluctuations of the MC templates can be determined from standard error propagation as
$\hat{\sigma}^{2}_{\rm MC}=\sum_{i=1}^{n}\left[\sum_{k=1}^{p}\left(\frac{\partial\hat{\mu}}{\partial B_{ik}}\right)^{2}(\Delta B_{ik})^{2}+\left(\sum_{k=1}^{p+1}{C}_{p+1\,k}B_{ik}\right)^{2}(\Delta f_{0,i})^{2}\right].$ (60)
@@
Equation (60) can be easily extended to generic covariances for ${\bf B}$ and ${\bf f}_{0}$. However, it should be noticed that the right-hand side of eq. (60) can only be evaluated using the MC prediction $\hat{\bf B}$ and $\hat{\bf f}_{0}$ derived from finite-size MC samples, so eq. (60) might be a biased estimator of the real variance of $\hat{\mu}$ under fluctuations of the MC templates. With this caveat in mind, we could still estimate the total variance of the maximum-likelihood estimator $\hat{\mu}$ under fluctuations of both the data and the MC samples as
$\hat{\sigma}_{\rm H,\,tot}\approx\sqrt{\hat{\sigma}_{\rm H}^{2}+\hat{\sigma}_{\rm MC}^{2}}.$ (61)
@@
Values of $\hat{\sigma}_{\rm MC}$ larger than the Barlow-Beeston lite expectation of $\hat{\sigma}_{\rm H}/\sqrt{k}$, or, equivalently values of $\hat{\sigma}_{\rm H,\,tot}$ in excess of the naive expectation $\hat{\sigma}_{\rm H}\sqrt{1+k^{-1}}$, should be taken as an indication that the lite approach is insufficient to provide a proper account of MC statistical uncertainties, suggesting that the full approach should be adopted instead. For example, for the toy study presented in section 2, we find that $\hat{\sigma}_{\rm MC}$ estimated from eq. (60) is a factor of $\sim 1.5$ larger than $\hat{\sigma}_{\rm H}/\sqrt{k}$ for the nominal model configuration. Interestingly, this ratio decreases monotonically towards $\sim 1$ when $N$ or $\epsilon$ are increased, or when $n$ is reduced, corroborating the observation that the Barlow-Beeston lite approach becomes more efficient in these limits. Conversely, we find that the ratio approaches a plateau of $\sim 1.6$ for $k\gg 1$, when $n$, $N$ and $\epsilon$ are fixed to their default values, implying that the Barlow-Beeston lite version will never be equivalent to its full implementation for at least the nominal model configuration [^10] In practice this difference is irrelevant when $k$ is large because the contribution from MC statistical uncertainties to the total variance will also be negligible in this limit.. A summary of numerical results is reported in table 4.

### v3

@@ 3.6 Asymptotic uncertainties from finite-size samples
The bias term generated by statistical fluctuations of the Jacobian matrix could be estimated in principle from eq. (3.2) by computing the covariance matrices of $\bf b$} and of the column vectors of $\bf A$}}, for example by bootstrap resampling, and using their nominal MC predictions $\hat{\bf b}$ and $\hat{\bf A}$ in place of their true, unknown values. The asymptotic Hessian uncertainty $\bar{\sigma}_{\rm H}={S}^{-1/2}$, i.e. the Hessian uncertainty in the limit of infinite MC statistics, could be determined from eq. (3.2) after subtracting the bias contribution.
@@
This result could be improved by taking the expectation values of $\hat{\sigma}_{{\rm H}}$ over several bootstrap re-samplings of the MC sample.
We can test the validity of eq. (48) for the toy model of section 2. Here, the mean Hessian uncertainty for $N=2\times 10^{6}$, $n=200$, and $\alpha=0.03$ is found to be $\hat{\sigma}_{{\rm H}}=0.0572$ ($0.0447$) when $k=1$ ($k=0.5$), from which the asymptotic uncertainty can be predicted to be $0.0956$ from eq. (48), in good agreement with the true value of $0.0950$. Extensions of eq (48) to the Barlow-Beeston lite likelihood are discussed in D.6.
@@
3.7 Testing the Barlow-Beeston lite approximation
@@
Suppose that $\bar{\sigma}_{\rm H}$ calculated from eq. (47) is in good agreement with $\hat{\sigma}_{\rm H}$, or, equivalently, that the Hessian uncertainty is stable when the size of the MC sample is reduced. One might still want to ensure that the lite version of the Barlow-Beeston method represents an adequate approximation for a given problem. For this purpose, we can use the formalism developed in the previous section. We start by noticing that the MLE estimators $\hat{\boldsymbol{\theta}}$ and $\hat{\mu}$, besides being functions of $\bf y$}, also depend on $\bf j$}, $\bf J$} and ${\bf f}_{0}$ through the matrix $\bf B$}, defined in D.1, and the vector $\bf d$} defined by eq. (30), which are all random variables of known variance. If the latter can be treated as independent and Gaussian-distributed variables with standard deviations $\Delta B_{ik}$ and $\Delta d_{i}$, the contribution to the variance of the MLE estimator $\hat{\mu}$ induced by fluctuations of the MC templates can be determined from standard error propagation as
$\hat{\sigma}^{2}_{\rm MC}=\sum_{i=1}^{n}\left[\sum_{k=1}^{p}\left(\frac{\partial\hat{\mu}}{\partial B_{ik}}\right)^{2}(\Delta B_{ik})^{2}+\left(\frac{\partial\hat{\mu}}{\partial d_{i}}\right)^{2}(\Delta d_{i})^{2}\right].$ (49)
@@
Analytical expressions for the two partial derivatives appearing in eq. (49) have been derived in D.7. Equation (49) can be easily extended to generic covariances for $\hat{\bf B}$ and $\hat{\bf d}$. However, we remark that the right-hand side of eq. (49) can only be evaluated using the MC prediction $\hat{\bf B}$ and $\hat{\bf d}$ derived from finite-size MC samples, so eq. (49) might be a biased estimator of the real variance of $\hat{\mu}$ under fluctuations of the MC templates. With this caveat in mind, we could still estimate the total variance of the maximum-likelihood estimator $\hat{\mu}$ under fluctuations of both the data and the MC samples as
$\hat{\sigma}_{\rm H}^{\rm BBfull}\approx\sqrt{\hat{\sigma}_{\rm H}^{2}+\hat{\sigma}_{\rm MC}^{2}}.$ (50)
@@
Values of $\hat{\sigma}_{\rm MC}$ larger than the Barlow-Beeston lite expectation of $\hat{\sigma}_{\rm H}/\sqrt{k}$, or, equivalently values of $\hat{\sigma}_{\rm H}^{\rm BBfull}$ in excess of the naive expectation $\hat{\sigma}_{\rm H}\sqrt{1+k^{-1}}$, should be taken as an indication that the lite approach is insufficient to provide a proper account of MC statistical uncertainties, suggesting that the full approach should be adopted instead. For example, for the toy study presented in section 2, we find that $\hat{\sigma}_{\rm MC}$ estimated from eq. (49) is a factor of $\sim 1.5$ larger than $\hat{\sigma}_{\rm H}/\sqrt{k}$ for the nominal model configuration. Interestingly, this ratio decreases monotonically towards $\sim 1$ when $N$ or $\epsilon$ are increased, or when $n$ is reduced, corroborating the observation that the Barlow-Beeston lite approach becomes more efficient in these limits. Conversely, we find that the ratio approaches a plateau of $\sim 1.6$ for $k\gg 1$, when $n$, NN and $\epsilon$ are fixed to their default values, implying that the Barlow-Beeston lite version will never be equivalent to its full implementation for at least the nominal model configuration [^11] In practice this difference is irrelevant when $k$ is large because the contribution from MC statistical uncertainties to the total variance will also be negligible in this limit.. A summary of numerical results is reported in table 6.
Finally, we remark that $\hat{\sigma}_{\rm MC}$ could also be estimated by considering the variance of $\hat{\mu}$ over multiple bootstrap re-samplings of the MC samples. However, eq. (49) offers the advantage that it is fully analytical and that, differently from resampling techniques, the estimators for $\bf B$} and $\bf d$} can be computed using the full statistical power of the simulated samples.